factor perfect-square trinomials and differences of squares.....
6Xsquared+2XY+Ysquared+3XY

Answer by checkley71(8403)

6X~2+2XY+Y~2+3XY=0 OR 6X~2+5XY+Y~2=0 FACTORS ARE(3X+Y)(2X+Y)=0 THUS X=-Y/3 &
X=-Y/2